# In-Context Learning Is Provably Bayesian Inference: A Generalization Theory for Meta-Learning

### v1

@@ 3.2 Posterior Variance: Inference-time Error
This theorem quantitatively justifies the empirical observation that ICL can quickly adapt to the specific task at hand, even when pretrained on a diverse mixture. Concretely, the posterior distribution over the task index, ${\mathcal{P}}_{I\mid D^{k}}$, concentrates exponentially fast on the true index $i^{\star}$ as $k$ grows. This result is consistent with empirical demonstrations. Panwar et al. (2024) show that in hierarchical mixtures, Transformers mimic the Bayes predictor based on the true task distribution. Also, the above theorem explains the “Bayesian scaling laws” of Arora et al. (2025), which model ICL’s error curves as repeated Bayesian updates, and under an ideal Bayesian learner, the task posterior converges to the true task as context grows.
@@
Compared to prior ICL theories, Theorem 3 can be seen as the general form of the result in Kim et al. (2024a), which showed that even when the function class used at pretraining is wider than the one at inference, the inference error depends only on the hardness of the latter class. Although Jeon et al. (2024) also mentions an irreducible error of ICL, our addition is to show that it manifests as Posterior Variance and that it approaches the minimax risk for the “true family” up to a small gap. Moreover, this phenomenon of ICL “selecting algorithms on the fly” is consistent with the theoretical results on in-context algorithm selection in generalized linear models and the Lasso (Akyürek et al., 2023; Bai et al., 2023; Zhang et al., 2024). Our result proves that even without assuming a specific algorithmic form, behavior close to optimal algorithm selection emerges through posterior concentration in mixture settings.
@@
The assumptions are fairly standard in the theory of sequential data and ensure that the in-context examples provide sufficient signal to rapidly rule out incorrect task types: (i) the supermartingale condition $\mathbb{E}[Z_{j,t}\mid{\mathcal{G}}_{t-1},I=i^{\star}]\leq-D_{j}<0$ (Williams, 1991) means each new observation, on average, decreases the predictive log-likelihood ratio of any wrong type $j$ against the true type; (ii) the Bernstein-type condition $\mathbb{E}\left[\exp\{\lambda(Z_{j,t}+D_{j})\}\mid{\mathcal{G}}_{t-1},I=i^{\star}\right]\leq\exp\left(\lambda^{2}\nu_{j}^{2}/2\right)$ (Bercu et al., 2015) yields the concentration of the cumulative log-likelihood ratio, so occasional misleading samples cannot outweigh the overall trend. Note that $D_{j}$ is the per-step information gap that favors the true task over the wrong type $j$, $\nu_{j}$ is the sub-exponential scale of the log-likelihood ratio, $b_{j}$ bounds the tail via the moment-generating-function (smaller means heavier tails), and $\min_{j\neq i^{\star}}D_{j}^{2}/8(\nu_{j}^{2}+b_{j}D_{j}/2)$ sets the uniform exponential rate at which posterior mass on wrong types decays with more context.
@@
In Appendix E, we consider a concrete regression problem (linear vs. series regression) and specify $\nu_{j},b_{j},D_{j}$ and $C$ that appear in Theorem 3. The results show that to make the task type identification error at most $\eta$, one requires the context length: $k\asymp\tfrac{\text{error variance}+\text{within-task variance}}{\text{true-task signal}}\log\tfrac{\text{\# of task types}}{\eta}$.
@@ 3.3 OOD Stability of the ICL risk
This result implies that the Bayes Gap is distributionally Lipschitz: its change across domains is controlled by (i) the smoothness $L$ of the Bayes predictor and (ii) the architectural regularity of $M_{\theta}$ through $L_{s}$, $L_{c}$, and $\mathrm{Lip}(\phi_{\theta})$. The penalty scales with the prompt-level Wasserstein shift and does not depend on the number of pretraining prompts $N$. In line with the findings of Zhang et al. (2024) that ICL is susceptible to input-distribution shifts, we find that only the Bayes Gap is affected by such shifts and quantify the magnitude of this effect.
For additional theories and detailed discussions, please see Appendix C.
@@ 4 Conclusion
Our analysis focuses on a uniform-attention Transformer, motivated by the permutation invariance of the Bayes predictor in our setup. Future work could explore how these results extend to more complex architectures with non-uniform attention, particularly in settings where sequential dependencies within the context are significant.
@@ Appendix A Notation and Definitions
  • $N^{\mathrm{seq}}_{2}(\alpha,{\mathcal{F}};z)$: The sequential covering number (Rakhlin et al., 2010; 2015) is the minimal size of a predictable $\alpha$-cover on a predictable tree $z$ with respect to $d_{2,\xi}(\cdot,\cdot;z)$ such that, for all $\xi\in\{\pm 1\}^{p}$ and all $f\in{\mathcal{F}}$, there exists $v$ in the cover with $d_{2,\xi}(f\circ z,v;z)\leq\alpha$.
@@
Appendix B Permutation Invariance and Justification for Uniform-Attention Transformers
@@
Hence, by convexity of the squared loss and conditional exchangeability (Lemma 1), permutation-averaging is a risk-reducing ensembling step. This is an instance of Rao–Blackwellization by group averaging under convex loss (Lehmann & Casella, 1998). Therefore, uniform-attention (mean-pooling) architectures are not only natural but also without loss of optimality in this setting.
@@
For further investigation of exchangeability of the Bayes predictor in the standard Bayesian statistics context, refer to Bernardo & Smith (1994); Gelman et al. (2013); Ghosal & van der Vaart (2017).
Appendix C Further Details on Out-of-Distribution Generalization
@@
Appendix D On the Hölder Condition of the Bayes Predictor
Although $y$ is not bounded in the prompt-generating process, the theorem imposes Hölder condition on bounded examples and queries. This is because if the noise follows a sub-Gaussian distribution, boundedness holds with high probability. Hence, the unbounded cases do not significantly affect the final result. Moreover, since the statement of the Theorem 2 is a bound on the expectation, it suffices that the Hölder condition holds with high probability.
@@ Appendix D On the Hölder Condition of the Bayes Predictor
  • Finite convex dictionary. $f_{a}=\sum_{j=1}^{J}a_{j}f^{(j)}$ with $\bm{a}\in\Delta^{J-1}$, each atom obeying $|f^{(j)}(\bm{x})|\leq B_{f}$ and $\|\nabla f^{(j)}(\bm{x})\|_{2}\leq L_{f}$ uniformly; take $\psi(\bm{x})=[f^{(1)}(\bm{x}),\ldots,f^{(J)}(\bm{x})]^{\top}$. Then $\bm{x}\mapsto f_{a}(\bm{x})$ is $L_{f}$-Lipschitz. (An example of a distribution on $\Delta^{J-1}$ is the logistic-normal distribution (Aitchison & Shen, 1980).)
@@
Appendix E Details of Theorem 3
@@
For conjugate normal models with bounded regressors $\phi,\psi$ and positive definite design covariances, standard ridge-risk bounds in series regression (§3.4 in van der Vaart & Wellner, 2023) give
@@
Appendix F Technical Lemmas
@@
Appendix G Proofs of the Main Results
@@
Step 3 (posterior concentration of the task index). Let $S_{j,k}:=\sum_{t=1}^{k}Z_{j,t}$ and $\lambda_{j,k}:=e^{S_{j,k}}$. By the assumption, $\{Z_{j,t}+D_{j}\}$ are conditionally sub-exponential supermartingale differences. Applying a Bernstein-type supermartingale inequality (Theorem 2.6 in Fan et al., 2015), for each $j\neq i^{\star}$,
@@
Appendix H Proofs of the Technical Lemmas
@@
Note that $\log N_{\infty,2}(\eta,\Phi;{\mathcal{N}})\leq\sum_{j=1}^{m}\log N_{\infty}\left(\tfrac{\eta}{\sqrt{m}},\Phi_{j};{\mathcal{N}}\right)\leq\sum_{j=1}^{m}\mathrm{Pdim}(\Phi_{j})\log\frac{C|{\mathcal{N}}|\sqrt{m}}{\eta}$. From Anthony & Bartlett (1999); Bartlett et al. (2019), using $\mathrm{Pdim}(\Phi)=\tilde{O}(m)$ for the coordinate-wise $[0,1]$-bounded ReLU features, the finite-set (size $|{\mathcal{N}}|$) covering bound gives
@@
Let $\bm{s}=\bm{n}/k$ and $\bm{t}=\bm{n}^{\prime}/k$ be points of $\Delta_{k}$. Construct an integer matrix $A=(A_{j\ell})$ with row sums $\bm{n}$ and column sums $\bm{n}^{\prime}$ (e.g., by the Northwest corner rule (Peyré & Cuturi, 2019)), and set $\pi:=A/k$. Then $\pi\in\Pi(s,t)$ is a feasible transport plan. Enumerating the $k$ pairs so that $(\bm{r}_{j(i)},\bm{r}_{\ell(i)})$ appears exactly $A_{j\ell}$ times, the Hölder condition yields
@@
We also note its piecewise linearity. By the Kantorovich–Rubinstein dual (Peyré & Cuturi, 2019) on a finite space,
@@
Next, fix $\bm{s}$ and vary $\bm{c},\bm{c}^{\prime}$. From the Hölder assumption on $M_{\mathrm{Bayes}}$ applied to $\bm{z}_{i}=(\bm{r}_{j(i)},\bm{c})$ and $\bm{z}^{\prime}_{i}=(\bm{r}_{j(i)},\bm{c}^{\prime})$ we obtain $|\rho_{\bm{c}}(\bm{v})-\rho_{\bm{c}^{\prime}}(\bm{v})|\leq L\|\bm{c}-\bm{c}^{\prime}\|_{2}^{\alpha}$ for all $\bm{v}\in\Delta_{k}$. By the McShane envelope (21), $(\bm{s},\bm{c})\mapsto\rho^{\star}_{\bm{c}}(\bm{s})$ is $\alpha$-Hölder in $\bm{c}$$|\rho^{\star}_{\bm{c}}(\bm{s})-\rho^{\star}_{\bm{c}^{\prime}}(\bm{s})|\leq L\|\bm{c}-\bm{c}^{\prime}\|_{2}^{\alpha}$. To meet the size of networks in Definition 2, we first apply a McShane-type $\alpha$-Hölder extension to the whole space $\mathbb{R}^{d_{\mathrm{feat}}}$, and then convolve only in the $\bm{c}$-direction with a standard mollifier (Appendix C.5 in Evans, 2010) $\eta_{h}$. This yields, for any $(\bm{s},\bm{c})$, $|\rho^{\star}_{\bm{c}}(\bm{s})-\rho^{\sharp}_{\bm{c}}(\bm{s})|\leq\left|\int\big(\rho^{\star}_{\bm{c}}(\bm{s})-\rho^{\star}_{{\bm{c}}-h{\bm{z}}}(\bm{s})\big)\,\eta(\bm{z})\,\mathrm{d}\bm{z}\right|\leq\int L\,\|h\bm{z}\|^{\alpha}\,\eta(\bm{z})\,\mathrm{d}\bm{z}\leq C_{\eta}\,L\,h^{\alpha}$, and $\mathrm{Lip}_{c}(\rho^{\sharp})\lesssim h^{\alpha-1}$ uniformly in $(\bm{s},\bm{c})$. In what follows we approximate $\rho^{\sharp}$ by a ReLU network and keep the same notation $\rho_{\theta}$.
(ii) ReLU approximation of the feature map. As in the current proof, each $\bm{u}\mapsto\phi_{j}(\bm{u})$ is $C^{\infty}$ on $0,1]^{d_{\mathrm{eff}}}$ with $\|\nabla\phi_{j}\|_{\infty}\lesssim\delta^{-1}$, hence by ReLU approximation (Yarotsky, 2017) there exists a ReLU network of depth $O(\log(1/\eta_{\phi}))$ and size $O(m\log(1/\eta_{\phi}))$ that uniformly approximates $\phi=(\phi_{1},\ldots,\phi_{m})$ on $\mathcal{U}$} with error $\eta_{\phi}\in(0,e^{-1})$. Additionally, we set spectral product
@@
(iii) ReLU approximation of the decoder. On the compact set $\Delta^{m-1}\times{\mathcal{C}}$, the map $(\bm{s},\bm{c})\mapsto\rho^{\sharp}_{\bm{c}}(\bm{s})$ is jointly Lipschitz with moduli $(L_{s},L_{c})$ from (i), and for each fixed $\bm{c}$ it is piecewise-linear in $\bm{s}$ (lower envelope of affine forms by the KR dual). Therefore, by standard approximation results for Lipschitz targets on a compact domain (Yarotsky, 2017), there exists a ReLU network $\rho_{\theta}:\Delta^{m-1}\times{\mathcal{C}}\to\mathbb{R}$ such that
@@
Recall that we work on standard Borel spaces (the Borel spaces associated with Polish spaces) so that regular conditional distributions (Durrett, 2019) exist. Accordingly, $\Pr(f\in\cdot\mid D^{k})$ and the quantities $\mathbb{E}[f(\bm{x}_{k+1})\mid D^{k}]$, $\mathrm{Var}(f(\bm{x}_{k+1})\mid D^{k})$ are well-defined.

### v2

@@ 3.2 Posterior Variance: Inference-time Error
This theorem quantitatively justifies the empirical observation that ICL can quickly adapt to the specific task at hand, even when pretrained on a diverse mixture. Concretely, the posterior distribution over the task index, ${\mathcal{P}}_{I\mid D^{k}}$, concentrates exponentially fast on the true index $i^{\star}$ as $k$ grows. This result is consistent with empirical demonstrations. panwar2024incontext show that in hierarchical mixtures, Transformers mimic the Bayes predictor based on the true task distribution. Also, the above theorem explains the “Bayesian scaling laws” of arora2025bayesian, which model ICL’s error curves as repeated Bayesian updates, and under an ideal Bayesian learner, the task posterior converges to the true task as context grows.
@@
Compared to prior ICL theories, Theorem 3 can be seen as the general form of the result in kim2024transformers, which showed that even when the function class used at pretraining is wider than the one at inference, the inference error depends only on the hardness of the latter class. Although jeon2024 also mentions an irreducible error of ICL, our addition is to show that it manifests as Posterior Variance and that it approaches the minimax risk for the “true family” up to a small gap. Moreover, this phenomenon of ICL “selecting algorithms on the fly” is consistent with the theoretical results on in-context algorithm selection in generalized linear models and the Lasso (akyurek2023what; bai2023transformers; zhang2024trained). Our result proves that even without assuming a specific algorithmic form, behavior close to optimal algorithm selection emerges through posterior concentration in mixture settings.
@@
The assumptions are fairly standard in the theory of sequential data and ensure that the in-context examples provide sufficient signal to rapidly rule out incorrect task types: (i) the supermartingale condition $\mathbb{E}[Z_{j,t}\mid{\mathcal{G}}_{t-1},I=i^{\star}]\leq-D_{j}<0$ (Williams_1991) means each new observation, on average, decreases the predictive log-likelihood ratio of any wrong type $j$ against the true type; (ii) the Bernstein-type condition $\mathbb{E}\left[\exp\{\lambda(Z_{j,t}+D_{j})\}\mid{\mathcal{G}}_{t-1},I=i^{\star}\right]\leq\exp\left(\lambda^{2}\nu_{j}^{2}/2\right)$ (BercuDelyonRio2015) yields the concentration of the cumulative log-likelihood ratio, so occasional misleading samples cannot outweigh the overall trend. Note that $D_{j}$ is the per-step information gap that favors the true task over the wrong type $j$, $\nu_{j}$ is the sub-exponential scale of the log-likelihood ratio, $b_{j}$ bounds the tail via the moment-generating-function (smaller means heavier tails), and $\min_{j\neq i^{\star}}D_{j}^{2}/8(\nu_{j}^{2}+b_{j}D_{j}/2)$ sets the uniform exponential rate at which posterior mass on wrong types decays with more context.
@@
In Appendix F, we consider a concrete regression problem (linear vs. series regression) and specify $\nu_{j},b_{j},D_{j}$ and $C$ that appear in Theorem 3. The results show that to make the task type identification error at most $\eta$, one requires the context length: $k\asymp\tfrac{\text{error variance}+\text{within-task variance}}{\text{true-task signal}}\log\tfrac{\text{\# of task types}}{\eta}$.
@@ 3.3 OOD Stability of the ICL risk
This result implies that the Bayes Gap is distributionally Lipschitz: its change across domains is controlled by (i) the smoothness $L$ of the Bayes predictor and (ii) the architectural regularity of $M_{\theta}$ through $L_{s}$, $L_{c}$, and $\mathrm{Lip}(\phi_{\theta})$. The penalty scales with the prompt-level Wasserstein shift and does not depend on the number of pretraining prompts $N$. In line with the findings of zhang2024trained that ICL is susceptible to input-distribution shifts, we find that only the Bayes Gap is affected by such shifts and quantify the magnitude of this effect.
For additional theories and detailed discussions, please see Appendix D.
@@ 4 Conclusion
Practical implications for model design and training.
For practitioners and model designers, our theory provides concrete guidance for optimizing in-context learning performance:
  1. -
    Architecture: Choosing the model’s feature dimension $m$ (hidden size) according to the scaling $m^{\star}\asymp(pN)^{\frac{d_{\mathrm{eff}}}{d_{\mathrm{eff}}+2\alpha}}$ is beneficial (Theorem 2). Moreover, when within-task examples are i.i.d., using permutation-invariant architectures such as uniform-attention (mean-pooling) Transformers, which match the symmetry of the Bayes-optimal predictor and avoid the curse of dimensionality in context length, is sufficient (cf. Appendix C).
  2. -
    Inference-time prompts: At inference time, an excessively long prompt may not be necessary, but the model does need at least a few examples to identify the task (about three in our experiments, cf. Figure 3). For more complex problems, the model needs enough examples to solve them, such that the minimax risk becomes sufficiently small (Theorem 3).
  3. -
    Distributional shift: One should match the input distribution at inference to that used in pretraining; our Wasserstein stability bound shows that distribution shifts primarily degrade the Bayes Gap (Theorem 4). When moderate domain shift is unavoidable, it is preferable to control the Lipschitz constants of the Transformer (e.g., via spectral normalization and output clipping) and then increase $p$ and $N$ and adjust $m$ according to the above scaling, rather than arbitrarily growing model size.
  4. -
    Benchmark design: Since ICL performance transitions from being dominated by task-type identification (small $k$) to within-family generalization (large $k$), few-shot benchmarks primarily measure task discrimination, while longer-context evaluations assess learning within the identified family.
@@
Our analysis focuses on a uniform-attention Transformer, motivated by the permutation invariance of the Bayes predictor in our setup. If inputs in a prompt are chosen adaptively (active learning, bandit-style data acquisition), non-uniform attention or explicitly sequential models would be beneficial. Future work could explore how these results extend to more complex architectures with non-uniform attention under dependent data settings. Our generalization analysis already uses tools from sequential learning, so the technical machinery can be compatible with order-dependent settings; what would change are the target Bayes predictor and its symmetry.
@@ Appendix A Notation and Definitions
  • $N^{\mathrm{seq}}_{2}(\alpha,{\mathcal{F}};z)$: The sequential covering number (Rakhlin2010online; Rakhlin2015) is the minimal size of a predictable $\alpha$-cover on a predictable tree $z$ with respect to $d_{2,\xi}(\cdot,\cdot;z)$ such that, for all $\xi\in\{\pm 1\}^{p}$ and all $f\in{\mathcal{F}}$, there exists $v$ in the cover with $d_{2,\xi}(f\circ z,v;z)\leq\alpha$.
@@
Appendix B Experiments
In this section, we empirically investigate (i) how the Bayes Gap of a Transformer depends on the number of pretraining prompts $N$ and their length $p$, and (ii) how the actual ICL prediction by the Transformer depends on the in-context length $k$.
Prompt Generation
We set the feature dimension $d_{\mathrm{feat}}=5$. Tasks are generated from a mixture of two regression families: Family 1 is linear regression $y=\bm{w}^{\top}\bm{x}+\varepsilon$, and Family 2 is regression via an element-wise nonlinear map $\phi(\bm{x})=\tanh(\bm{x})$, i.e., $y=\bm{w}^{\top}\phi(\bm{x})+\varepsilon$. The mixture weights are $(\alpha_{1},\alpha_{2})=(0.5,0.5)$. The weight vector $\bm{w}$ for each family is drawn from $N(\bm{\mu}_{i},\tau^{2}I)$ with $\tau=1$, $\bm{\mu}_{1}=(3,3,3,3,3)$ and $\bm{\mu}_{2}=(-3,-3,-3,-3,-3)$. The observation noise $\varepsilon$ is zero-mean Gaussian with $\sigma_{\varepsilon}=0.1$. This process generates prompts $P^{k}=(\bm{x}_{1},y_{1},\,\ldots,\,\bm{x}_{k},y_{k},\,\bm{x}_{k+1})\in(\mathbb{R}^{d_{\mathrm{feat}}}\times\mathbb{R})^{k}\times\mathbb{R}^{d_{\mathrm{feat}}}$ and the target $y_{k+1}\in\mathbb{R}$.
Architecture
As in oko2024pretrained and garg2022case [^2] https://github.com/dtsip/in-context-learning, MIT License., we adopt a GPT-2 model (radford2019language; wolf-etal-2020-transformers), with 12 layers, 8 heads, a hidden size of 256, and positional embeddings removed. The input is a token sequence of length $2k+1$: $P^{k}=(\bm{x}_{1},y_{1},\dots,\bm{x}_{k},y_{k},\bm{x}_{k+1})$.
Evaluation of Bayes Gap
For any prompt $(D^{k},\bm{x}_{k+1})$, since each family admits a conjugate Bayesian update, we can compute the mixture posterior $\pi_{i}(D^{k})$ and the Bayes predictor $M_{\mathrm{Bayes}}$ in closed form (gelman2013bda3). At inference time, using token sequences of length $2p+1$, we compute the difference between the model’s prediction and the Bayes prediction $M_{\mathrm{Bayes}}(D^{k},\bm{x}_{k+1})$. We average this over 1000 trials and report the mean as the Bayes Gap.
  • NN-Sweep (fixed $p$): For $p\in\{5,10,15\}$, we train the model on data with $N\in\{128,256,\dots,8192\}$ and report the Bayes Gap.
  • $p$-Sweep (fixed $N$): For $N\in\{500,1000,2000\}$, we vary $p\in\{1,2,4,6,8,10,12,14\}$ and report the Bayes Gap.
During pretraining, we perform the ERM as in (6) using Adam (kingma2015adam; paszke2019pytorch) with a learning rate of $0.00005$, a batch size of 64, and 10000 steps.
Figure 2: Behavior of the Bayes Gap (left: NN-sweep, right: $p$-sweep). The left panel fixes $p\in\{5,10,15\}$ and varies the number of pretraining prompts NN; the right panel fixes $N\in\{500,1000,2000\}$ and varies the context length $p$ in pretraining. In both cases, the Bayes Gap decreases generally as NN or $p$ increases, demonstrating that longer contexts and more pretraining improve approximation to the Bayes predictor.
Figure 2 illustrates how the Bayes Gap depends jointly on the number of pretraining prompts $N$ and the context length $p$. In the $N$-sweep (left panel), larger $p$ values start with smaller gaps and maintain this advantage as $N$ grows. In the $p$-sweep (right panel), increasing $p$ steadily reduces the gap for any fixed $N$, with higher $N$ curves lying lower overall.
Evaluation of In-Context Error
We pretrain the model on a dataset with $N=12800000$ and $p=15$ using Adam, which involves 200000 steps of online learning, sampling 64 data points per step, similar to (garg2022case). On $1000$ test prompts, the Bayes Gap is below $10^{-2}$, so the pretrained Transformer reasonably approximates $M_{\mathrm{Bayes}}$ (cf. Theorem 2). We then investigate: (i) the prediction accuracy of the Transformer via the mean squared error (MSE) with respect to $y$, and (ii) Bayes-like behavior of the Transformer via linear probing: generating inputs, recovering the regression coefficients implied by the Transformer’s outputs (inverting the induced linear map), and comparing against the optimal Bayesian coefficients. The probed coefficients are averaged over $2000$ independent input–output sets.
Figure 3: In-context error under task mixtures (left: predictive MSE; right: parameter-estimation MSE). As the context length $k$ increases, both predictive error for the next label $y_{k+1}$ from $P^{k}$ (left) and parameter-estimation error (right) decrease monotonically. The Transformer closely tracks the mixture Bayes predictor and, with sufficient context, approaches the oracle Bayes curve that knows the true task family. This demonstrates the rapid concentration of the task-index posterior under growing context and the corresponding shrinkage of the irreducible term.
Figure 3 illustrates the inference-time behavior of the Transformer. Note that Bayes (mixture) corresponds to the Bayes-optimal in-context predictor $M_{\mathrm{Bayes}}$ that conditions on the mixture prior over task families, whereas Bayes (oracle) denotes the Bayes predictor that knows the true task family at inference time. First, with sufficiently large pretraining, the Transformer tracks the Bayes (mixture) curve almost exactly and its error decays toward zero, consistent with our theory (Theorem 2). In terms of parameter estimation, the Transformer likewise reproduces the Bayes-mixture behavior, supporting the view that it performs Bayesian inference in context. Moreover, the gap between the Transformer and Bayes (oracle) diminishes rapidly as the number of in-context examples $k$ increases and essentially vanishes around $k\approx 3\text{--}4$. This indicates that the task type identification error quickly diminishes and the remaining bottleneck is the intrinsic (optimal) error of the true task family (Theorem 3).
Appendix C Permutation Invariance and Justification for Uniform-Attention Transformers
@@
Hence, by convexity of the squared loss and conditional exchangeability (Lemma 1), permutation-averaging is a risk-reducing ensembling step. This is an instance of Rao–Blackwellization by group averaging under convex loss (LehmannCasella1998). Therefore, uniform-attention (mean-pooling) architectures are not only natural but also without loss of optimality in this setting.
@@
For further investigation of exchangeability of the Bayes predictor in the standard Bayesian statistics context, refer to BernardoSmith1994; gelman2013bda3; ghosal2017fundamentals.
Appendix D Further Details on Out-of-Distribution Generalization
@@
Appendix E On the Hölder Condition of the Bayes Predictor
Although $y$ is not bounded in the prompt-generating process, the theorem imposes the Hölder condition on bounded examples and queries. This is because if the noise follows a sub-Gaussian distribution, boundedness holds with high probability. Hence, the unbounded cases do not significantly affect the final result. Moreover, since the statement of Theorem 2 is a bound on the expectation, it suffices that the Hölder condition holds with high probability.
@@ Appendix E On the Hölder Condition of the Bayes Predictor
  • Finite convex dictionary. $f_{a}=\sum_{j=1}^{J}a_{j}f^{(j)}$ with $\bm{a}\in\Delta^{J-1}$, each atom obeying $|f^{(j)}(\bm{x})|\leq B_{f}$ and $\|\nabla f^{(j)}(\bm{x})\|_{2}\leq L_{f}$ uniformly; take $\psi(\bm{x})=[f^{(1)}(\bm{x}),\ldots,f^{(J)}(\bm{x})]^{\top}$. Then $\bm{x}\mapsto f_{a}(\bm{x})$ is $L_{f}$-Lipschitz. (An example of a distribution on $\Delta^{J-1}$ is the logistic-normal distribution (AitchisonShen1980).)
@@
Appendix F Details of Theorem 3
@@
For conjugate normal models with bounded regressors $\phi,\psi$ and positive definite design covariances, standard ridge-risk bounds in series regression (§3.4 in vaart_wellner_weak_2023) give
@@
Appendix G Technical Lemmas
@@
Appendix H Proofs of the Main Results
@@
Step 3 (posterior concentration of the task index). Let $S_{j,k}:=\sum_{t=1}^{k}Z_{j,t}$ and $\lambda_{j,k}:=e^{S_{j,k}}$. By the assumption, $\{Z_{j,t}+D_{j}\}$ are conditionally sub-exponential supermartingale differences. Applying a Bernstein-type supermartingale inequality (Theorem 2.6 in FanGramaLiu2015), for each $j\neq i^{\star}$,
@@
Appendix I Proofs of the Technical Lemmas
@@
Note that $\log N_{\infty,2}(\eta,\Phi;{\mathcal{N}})\leq\sum_{j=1}^{m}\log N_{\infty}\left(\tfrac{\eta}{\sqrt{m}},\Phi_{j};{\mathcal{N}}\right)\leq\sum_{j=1}^{m}\mathrm{Pdim}(\Phi_{j})\log\frac{C|{\mathcal{N}}|\sqrt{m}}{\eta}$. From AnthonyBartlett1999; Bartlett2019nearly, using $\mathrm{Pdim}(\Phi)=\tilde{O}(m)$ for the coordinate-wise $[0,1]$-bounded ReLU features, the finite-set (size $|{\mathcal{N}}|$) covering bound gives
@@
Let $\bm{s}=\bm{n}/k$ and $\bm{t}=\bm{n}^{\prime}/k$ be points of $\Delta_{k}$. Construct an integer matrix $A=(A_{j\ell})$ with row sums $\bm{n}$ and column sums $\bm{n}^{\prime}$ (e.g., by the Northwest corner rule (peyré2020computational)), and set $\pi:=A/k$. Then $\pi\in\Pi(s,t)$ is a feasible transport plan. Enumerating the $k$ pairs so that $(\bm{r}_{j(i)},\bm{r}_{\ell(i)})$ appears exactly $A_{j\ell}$ times, the Hölder condition yields
@@
We also note its piecewise linearity. By the Kantorovich–Rubinstein dual (peyré2020computational) on a finite space,
@@
Next, fix $\bm{s}$ and vary $\bm{c},\bm{c}^{\prime}$. From the Hölder assumption on $M_{\mathrm{Bayes}}$ applied to $\bm{z}_{i}=(\bm{r}_{j(i)},\bm{c})$ and $\bm{z}^{\prime}_{i}=(\bm{r}_{j(i)},\bm{c}^{\prime})$ we obtain $|\rho_{\bm{c}}(\bm{v})-\rho_{\bm{c}^{\prime}}(\bm{v})|\leq L\|\bm{c}-\bm{c}^{\prime}\|_{2}^{\alpha}$ for all $\bm{v}\in\Delta_{k}$. By the McShane envelope (21), $(\bm{s},\bm{c})\mapsto\rho^{\star}_{\bm{c}}(\bm{s})$ is $\alpha$-Hölder in $\bm{c}$$|\rho^{\star}_{\bm{c}}(\bm{s})-\rho^{\star}_{\bm{c}^{\prime}}(\bm{s})|\leq L\|\bm{c}-\bm{c}^{\prime}\|_{2}^{\alpha}$. To meet the size of networks in Definition 2, we first apply a McShane-type $\alpha$-Hölder extension to the whole space $\mathbb{R}^{d_{\mathrm{feat}}}$, and then convolve only in the $\bm{c}$-direction with a standard mollifier (Appendix C.5 in Evans2010PDE) $\eta_{h}$. This yields, for any $(\bm{s},\bm{c})$, $|\rho^{\star}_{\bm{c}}(\bm{s})-\rho^{\sharp}_{\bm{c}}(\bm{s})|\leq\left|\int\big(\rho^{\star}_{\bm{c}}(\bm{s})-\rho^{\star}_{{\bm{c}}-h{\bm{z}}}(\bm{s})\big)\,\eta(\bm{z})\,\mathrm{d}\bm{z}\right|\leq\int L\,\|h\bm{z}\|^{\alpha}\,\eta(\bm{z})\,\mathrm{d}\bm{z}\leq C_{\eta}\,L\,h^{\alpha}$, and $\mathrm{Lip}_{c}(\rho^{\sharp})\lesssim h^{\alpha-1}$ uniformly in $(\bm{s},\bm{c})$. In what follows we approximate $\rho^{\sharp}$ by a ReLU network and keep the same notation $\rho_{\theta}$.
(ii) ReLU approximation of the feature map. As in the current proof, each $\bm{u}\mapsto\phi_{j}(\bm{u})$ is $C^{\infty}$ on $0,1]^{d_{\mathrm{eff}}}$ with $\|\nabla\phi_{j}\|_{\infty}\lesssim\delta^{-1}$, hence by ReLU approximation (Yarotsky2017error) there exists a ReLU network of depth $O(\log(1/\eta_{\phi}))$ and size $O(m\log(1/\eta_{\phi}))$ that uniformly approximates $\phi=(\phi_{1},\ldots,\phi_{m})$ on $\mathcal{U}$} with error $\eta_{\phi}\in(0,e^{-1})$. Additionally, we set spectral product
@@
(iii) ReLU approximation of the decoder. On the compact set $\Delta^{m-1}\times{\mathcal{C}}$, the map $(\bm{s},\bm{c})\mapsto\rho^{\sharp}_{\bm{c}}(\bm{s})$ is jointly Lipschitz with moduli $(L_{s},L_{c})$ from (i), and for each fixed $\bm{c}$ it is piecewise-linear in $\bm{s}$ (lower envelope of affine forms by the KR dual). Therefore, by standard approximation results for Lipschitz targets on a compact domain (Yarotsky2017error), there exists a ReLU network $\rho_{\theta}:\Delta^{m-1}\times{\mathcal{C}}\to\mathbb{R}$ such that
@@
Recall that we work on standard Borel spaces (the Borel spaces associated with Polish spaces) so that regular conditional distributions (Durrett_2019) exist. Accordingly, $\Pr(f\in\cdot\mid D^{k})$ and the quantities $\mathbb{E}[f(\bm{x}_{k+1})\mid D^{k}]$, $\mathrm{Var}(f(\bm{x}_{k+1})\mid D^{k})$ are well-defined.